3) Determine the tension in each segment of the cable shown below. A E Чр B 15m р B C YKN SKM 18 КМ .am 1-3m 3. . •3 ouTThis is all the information given in the problem. There is a solution of a similar problem in the notes for the class and the values are just changed.
Answer #1

Draw the free Body diagram. Ay Ax & Ex 8 b Ey 1 4 Чk, 5KN BKN EME=0 Take moment about E, ECA Ayx 17 ) +(5X.15) +(18x 127 + HTalce moment about c, IM = 0 (Aq X 155 ) + (5*3 )-(Ayx 55 )=0 15 Ant 15-89.1 da = 4.94 KN Druce the face Body diagrama Ag. 1Take momnt abouto, IMO=0 [(4,4 40 9b) +(6x 12)+(x 91 Ayx 14 [(4.qux y ) +(8x12) +(1869)={17.82x143)=0 Yo 5.562 At Joint A qtyFrom figure-2, fand the Tergica in cable BC. the angle at point B. fan OB = 15-4B 15-74214 3 3 Qg = 68.930 Rescle the forceshorizontal Resolve the force along direction, & Fr=0. - Aut To con op = 0 -4.944 Tro COS(46.360) =0 Tco=70158 KN From figure

I have provided conceptually correct answer.

If you have any doubts please ask me.

Please provide positive rating and thank you.